# Manuscript

### v1

@@
Manuscript
Version One
ON THE REALIZABILITY OF
PRIME CONJECTURES
IN HEYTING ARITHMETIC
Milan Rosko
(November 2025)
@@
Keywords: Proof Theory, Constructive Logic, Realizability;
Heyting Arithmetic; Goldbach’s Conjecture; Geometric Logic
MSC: 03F55, 03F03, 11A41, 68Q05, 11U05
Milan Rosko is from University of Hagen, Germany
Email: Q1012878 @@ studium.fernuni-hagen.de
ORCID: 0009-0003-1363-7158
Licensed under “Deed” ΓΔ  – creativecommons.org/licenses/by/4.0
@@ 1 Introduction
Let mathematics correspond to the greatest rigor after assumption; then logic will be the rigor before assumption. We examine the boundary where the two intersect. Classical arithmetic proceeds as though its objects—numbers, primes, and sums—are already well formed, whereas logic asks whether the rules that make such objects definable can themselves support constructive realization. The contrast becomes evident in statements such as Goldbach’s:
@@
Our analysis shows that any uniform construction transforming an even input into explicit prime summands would conflict with the normalization theorems of constructive proof theory rather than with any property of the primes themselves. It concerns the structure of realizability in arithmetic, not the distribution of prime numbers.
The distinction between additive and multiplicative reasoning marks a precise logical frontier. To demonstrate “compositeness” is to exhibit explicit factors, a $\Sigma_{1}$ act. To demonstrate “primality” is to assert the universal absence of nontrivial factors, a $\Pi_{1}$ claim.
@@ 2.1 Semantics
Proofs correspond to explicit constructions (realizers):
@@ 2.2 Thesis
Example 2.3.
Recall the Arithmetic Hierarchy of Kleene [1943]. It classifies subsets of $\mathbb{N}$ and arithmetical formulas according to the alternation of quantifiers in their definitions.
@@
Thesis 2.4 (Realizability Barrier).
@@ 2.4 Constructive Translation
This statement combines $\Sigma_{1}$ witnesses $(p,q)$ with $\Pi_{1}$ side-conditions $(\operatorname{Prime_{\Pi}}(p),\operatorname{Prime_{\Pi}}(q))$. A constructive proof of (2.8) in HA would yield, via the Curry–Howard Correspondence, a total primitive-recursive operator
$E:\mathbb{N}\longrightarrow\mathbb{N}^{5},\quad E(n)=\langle p,q,a,b,c\rangle,$ (2.9)
where $a,b$ realize the predicates $\operatorname{Prime_{\Pi}}(p),\operatorname{Prime_{\Pi}}(q)$, and $c$ realizes $p+q=n$. Such an operator would act as a uniform $\Sigma_{1}$ extractor for $\Pi_{1}$ evidence, contradicting the strong normalization of HA proofs, see [Kreisel, 1951].
@@ 5.1 Inductive Inference and Semantic Reduction
We simulate classical logic within our constructive framework by inductive reasoning, examining structural correspondences. We appeal to the Theory of Inductive Inference, see Solomonoff [1964a, b] and to the Church–Turing Thesis: any realizible arithmetical predicate can be represented by a minimal constructive schema carrying equivalent realization. For primality,
@@ 5.4 Euler Product from Fixed-Point
Its fixed point is the constant $\Phi$, which quantifies the equilibrium between constructive expansion ($\Sigma_{1}$ search) and reflective contraction ($\Pi_{1}$ verification). See Section Context for related discussion.
@@ 5.5 Decomposition and Factorization
Dirichlet addition is ordinary summation, and multiplication acts additively on exponents, see [Apostol, 1976]:
@@
where $\Lambda$ denotes the Von Mangoldt Function, see [von Mangoldt, 1895].
@@ 7.2 Primes as Proof of Knowledge
For example, Quantum computation accelerates instance evaluation but does not transform a given nonconstructive predicate into a realization. Logical non-uniformity persists across such computation, see [Abramsky, 2004].
@@ 8 Discussion
We close with the central idea: the resolution of future obstructions in mathematics, logic, physics, and computation will not emerge from the isolated treatment of individual problems as dichotomies, but from their systematic integration under a unified arithmetic oracle—an architecture capable of identifying, rather than eliminating, the shared constructive limits that underlie them all:
@@ Acknowledgments
Context
For discussion of $\Phi$-adic” recursion, see arXiv:2510.08934 or Rosko [2025].
Final Remarks
The author welcomes scholarly correspondence and constructive dialogue. No conflicts of interest are declared. This research received no external funding.

### v2

# ON THE REALIZABILITY OF PRIME CONJECTURES IN HEYTING ARITHMETIC

@@
Abstract
@@
Subject: Realizability, Constructive Logic, Primes
Proof Theory, Recursive Arithmetic, Geometric Logic
@@ 1 Introduction
Let mathematics correspond to the greatest rigor after assumption; then logic will be the rigor before assumption. In this spirit, we examine the boundary where the two intersect. Classical arithmetic proceeds as though its objects—numbers, primes, and sums—are already well formed, whereas logic asks whether the rules that make such objects definable can themselves support constructive realization. The contrast becomes evident in statements such as Goldbach’s:
@@
Our analysis shows that any uniform construction transforming an even input into explicit prime summands would conflict with the normalization theorems of constructive proof theory rather than with any property of the primes themselves. It concerns the structure of realizability in arithmetic, not the distribution of prime numbers. The distinction between additive and multiplicative reasoning marks a precise logical frontier. To demonstrate “compositeness” is to exhibit explicit factors, a $\Sigma_{1}$ act. To demonstrate “primality” is to assert the universal absence of nontrivial factors, a $\Pi_{1}$ claim.
@@ 2.1 Semantics
Proofs correspond to realizers:
@@ 2.2 Thesis
Ansatz 2.3 (BHK Witnessability).
Throughout, existence is understood in the BHK sense: a statement of the form
$\forall x\,(\varphi(x)\Rightarrow\exists y\,\psi(x,y))$ (2.3)
is treated not as a mere classical truth condition, but as a specification of a uniform constructive operator. Concretely, a proof in HA of such a sentence is taken to determine a total functional $t$ such that
$\forall x\,(\varphi(x)\Rightarrow\psi(x,t(x))).$ (2.4)
Accordingly, we do not regard the possibility that witnesses might exist in principle without being uniformly producible as a semantic resource. In this work, arithmetical existence claims are evaluated only insofar as they admit explicit constructive realization under normalization. Consider The Teapot of Russell [1952] as an apt anlogy.
Ansatz 2.4 (Arithmetical Hierarchy).
Recall the Kleene–Mostowski Hierarchy by Kleene [1943]. It classifies subsets of $\mathbb{N}$ and arithmetical formulas according to the alternation of quantifiers in their definitions.
@@
Thesis 2.5 (Realizability Barrier).
@@
Thesis 2.6 (Uniform Realizability Barrier).
No closed term $t$ in any strongly normalizing realizability calculus adequate for HA (e.g. system $\mathcal{T}$ or an equivalent typed $\lambda$-calculus with primitive recursion) can uniformly transform $\Pi_{1}$ evidence into explicit $\Sigma_{1}$ witnesses in the following sense: there is no total realizer $t$ such that, for all $n$,
$t(n)\;\Vdash\;\Bigl(\Pi_{1}(n)\Rightarrow\exists y\,\Sigma_{1}(n,y)\Bigr)\quad\Longrightarrow\quad\exists y\,\bigl(t(n)=y\wedge\Sigma_{1}(n,y)\bigr).$ (2.9)
Equivalently, $\Pi_{1}$ predicates do not admit uniform $\Sigma_{1}$-Skolemization within the normalization regime of constructive proof theory.
@@ 2.4 Constructive Translation
Definition 2.9 (Base Calculus).
Instead of presenting constructive arithmetic solely as a first-order theory, one may work in a typed $\lambda$-calculus $\Lambda_{\mathbb{N}}$ with natural numbers, primitive recursion, and inductive types, whose equational theory corresponds to HA under the Curry–Howard correspondence. We call a functional constructively definable if it is represented by a closed term of $\Lambda_{\mathbb{N}}$; equivalently, if it is definable in Gödel’s system $\mathcal{T}$. All extracted witness mechanisms discussed below are to be understood in this $\lambda$-definability sense.
Observation.
A proof of (2.11) in HA would, under any standard realizability or functional interpretation, induce a uniform closed term $t$ in a strongly normalizing typed $\lambda$-calculus interpreting HA, such that for each even input $n$ the term $t(n)$ computes explicit numerals $\langle p,q\rangle$ satisfying $p+q=n$ together with realizers of the side conditions $\operatorname{Prime_{\Pi}}(p)$ and $\operatorname{Prime_{\Pi}}(q)$. In other words, a proof of (2.11) would not merely assert the existence of prime summands, but would entail a uniform constructive mechanism producing them, together with the global $\Pi_{1}$ evidence that they are prime. It is precisely this uniformity, rather than any arithmetical property of primes, that contradicts the non-uniformization theorems of constructive proof theory.
Exposition.
One might ask: if a uniform witness extractor is unavailable in HA, then in which system could it exist? The point is that any such extractor amounts to a global choice principle for $\Pi_{1}$-premised existence, and thus requires principles that exceed normalization-governed constructive arithmetic. Typical extensions that admit such uniformization include: (i) classical logic (via $\neg\neg$-elimination and Skolemization), (ii) choice principles for arithmetical predicates, or (iii) non-normalizing computational interpretations (e.g. general recursion or bar recursion). Our barrier is therefore not contingent on HA syntactically: it persists across all systems whose proof semantics remain strongly normalizing and computationally sound for arithmetic.
@@ 5.1 Inductive Inference and Semantic Reduction
We simulate classical logic within our constructive framework by inductive reasoning, examining structural correspondences. We appeal to the Theory of Inductive Inference, cf. Solomonoff [1964a, b] and to the Church–Turing Thesis: any realizible arithmetical predicate can be represented by a minimal constructive schema carrying equivalent realization. For primality,
@@ 5.4 Euler Product from Fixed-Point
Its fixed point is the constant $\Phi$, which quantifies the equilibrium between constructive expansion ($\Sigma_{1}$ search) and reflective contraction ($\Pi_{1}$ verification), see Section Acknowledgments for related discussion.
@@ 5.5 Decomposition and Factorization
Dirichlet addition is ordinary summation, and multiplication acts additively on exponents, cf. [Apostol, 1976]:
@@
where $\Lambda$ denotes the Von Mangoldt Function, cf. [von Mangoldt, 1895].
@@ 7.2 Primes as Proof of Knowledge
For example, Quantum computation accelerates instance evaluation but does not transform a given nonconstructive predicate into a realization. Logical non-uniformity persists across such computation, cf. [Abramsky, 2004].
@@ 8 Discussion
We close with the central idea: the resolution of future obstructions in mathematics, logic, physics, and computation will not emerge from the isolated treatment of individual problems as dichotomies, but from their systematic integration under a general arithmetic oracle—an architecture capable of identifying, rather than eliminating, the shared constructive limits that underlie obstructions: